# Combining Deterministic Enhanced Conditions with Dual-Streaming Encoding for Diffusion-Based Speech Enhancement

### v1

@@ II-B Score-based Diffusion Model for Speech Enhancement
where $z\sim\mathcal{N_{C}}(\mathrm{z};0,\mathbf{I})$ is a sampled white Gaussian noise with an identity matrix $\mathbf{I}$.
III Proposed Method
We propose combining the deterministic and diffusion models to meet the partially stochastic function property of the supervised SE. It contains a deterministic model and a diffusion model as shown in Fig. 1. In the deterministic model, we propose a granularity progressive deterministic model COFFEE by leveraging the advantages of both high speed down-sampling processing and high performance fine granularity processing as shown in Fig. 1. The bottleneck embedding, reconstructed features in the decoder, and the T-F bins are enhanced with fine-granularity processing gradually.
To better integrate the deterministic and diffusion models, we propose a Repair-Diffusion model for SE (RDM-SE). We design the encoder of RDM-SE based on the characteristics of deterministic and diffusion processes. Within the encoder, the features are processed through several deterministic-diffusion layers and diffusion-specific layers. Deterministic information is explicitly introduced into the shared encoder layers by computing the deterministic loss. The deterministic-diffusion shared layers perform downsampling only along the frequency axis to retain sufficient information for deterministic information repair. The diffusion-specific layers focus on extracting embeddings only for diffusion. Besides, the deterministic models affect the choice between using deterministic-only and deterministic-noisy conditions. Thus, we further improve the RDM-SE with dual-stream encoding to simultaneously utilize both deterministic-only and deterministic-noisy conditions.
Figure 1: Flowchart of the proposed method. It includes the proposed COFFEE deterministic model and the proposed RDM-SE. Within COFFEE, we introduce Time-Embedding (T-E) bin-level, Time-Reconstructed Feature (T-RF) bin-level, and Time-Frequency (T-F) bin-level processing. Within RDM-SE, the skip connections between the encoder and decoder are omitted in the figure for clarity.
III-A Deterministic Method: COFFEE
It contains an encoder, a decoder, and the bottleneck processing layers. The encoder reduces the resolution of the input feature and extracts high-level representations. The encoder layer processes the input complex feature $E=E_{r}+jE_{i}$ with complex Conv2d. $E$, $E_{r}$, and $E_{i}$ represent the complex, real, and imaginary parts of the complex spectrogram or encoded features. $j$ is the imaginary unit as $j^{2}=-1$. The complex-valued filter $W=W_{r}+jW_{i}$ is used for convolution operation:
$E_{out}=(E_{r}*W_{r}^{e}-E_{i}*W_{i}^{e})+j(E_{r}*W_{i}^{e}+E_{i}*W_{r}^{e}),$ (9)
the real-valued matrices $W_{r}^{e}$ and $W_{i}^{e}$ represent the real and imaginary parts of the complex convolution kernel, respectively. $W_{r}^{e}$ and $W_{i}^{e}$ are implemented with the Conv2d layer with different parameters. Batch normalization and PReLU follow the Conv2d layers. $E_{out}$ is the filtered feature.
@@ III-A Deterministic Method: COFFEE
where the $W^{b}$ is implemented with the T-E processing block. Fig. 1 shows the structure of the T-E processing Block. It contains an Intra-LSTM, an Inter-LSTM, and an attention mechanism to enhance full-band and sub-band modeling capabilities. Two additional linear layers with complex convolution are included for the separate modeling. The Intra-LSTM, Inter-LSTM, and Attention are similar to the process of processing features described in Section II-A. The T-E bin level feature is processed with a large dimension reduction compared with the original T-F spectrogram.
@@
Finally, after the feature is upsampled to the same resolution as the original noisy feature, a T-F processing Block is used to make the final refining. The T-F processing block has the same structure as the T-RF processing block.
@@
$CRM=\frac{Y_{r}S_{r}+Y_{i}S_{i}}{Y_{r}^{2}+Y_{i}^{2}}+j\frac{Y_{r}S_{i}-Y_{i}S_{r}}{Y_{r}^{2}+Y_{i}^{2}}$ (12)
where $Y_{r}$ and $Y_{i}$ represent the real and imaginary parts of the noisy complex spectrogram, respectively. The $S_{r}$ and $S_{i}$ represent the real and imaginary parts of the clean complex spectrogram, respectively. According to the polar coordinates of Cartesian coordinate representation $\tilde{M}=\tilde{M}_{r}+j\tilde{M}_{i}$:
$\tilde{M}_{\text{mag}}&=\sqrt{\tilde{M}_{r}^{2}+\tilde{M}_{i}^{2}},\\ \tilde{M}_{\text{phase}}&=\arctan 2(\tilde{M}_{i},\tilde{M}_{r}).\end{aligned}\right.$ (13)
@@
$\tilde{S}=Y_{mag}\cdot\tilde{M}_{mag}\cdot e^{(Y_{phase}+\tilde{M}_{phase})}$ (14)
where $Y_{mag}$ and $Y_{phase}$ represent the magnitude and phase information of the noisy speech. The training loss is based on the MSE between clean and enhanced spectrogram:
$\mathcal{L}_{Deter}=|\tilde{S}-S|^{2},$ (15)
where $S$ represents the clean spectrogram.
III-B Diffusion Model: RDM-SE
@@
[…]
@@
Figure 2: Flowchart of the proposed dual-streaming encoding RDM-SE (DERDM-SE). The skip connections between the encoder and decoder are omitted in the figure for clarity.
We propose Repair-Diffusion to better utilize deterministic information for diffusion (RDM-SE). It is based on dual-decoder sturcture. The model contains an encoder, a bottleneck processing module, a diffusion decoder, and a deterministic repair decoder. Within the encoder, the feature is processed by several deterministic-diffusion shared layers and diffusion-specific layers. The conditions and diffused features are first processed by the encoder:
$\begin{cases}\mathbf{e^{\prime}_{dd}}=\text{Enc}_{\text{DD}}(\mathbf{x^{\prime}}_{t},\mathbf{x^{\prime}},\mathbf{y},t)\\ \mathbf{e^{\prime}_{ds}}=\text{Enc}_{\text{DS}}(\mathbf{e^{\prime}_{dd}},t)\end{cases}$ (21)
where $\text{Enc}_{\text{DD}}$ and $\text{Enc}_{\text{DS}}$ represent the deterministic-diffusion shared layers and diffusion-specific layers, respectively. $\mathbf{e^{\prime}_{dd}}$ and $\mathbf{e^{\prime}_{ds}}$ denote the deterministic-diffusion shared embeddings and diffusion-specific embeddings, respectively. The deterministic-diffusion shared layer consists of several ResNet blocks and a T-RF bin processing block. It performs downsampling only along the frequency axis to preserve the repair of deterministic information. The deterministic-specific layer consists of several ResNet blocks. The attention layer is also applied to some layers but not all. It performs downsampling either along the time axis or the time-frequency axis.
@@
It should be noted that not all encoder layers were used for the repair of deterministic information. As a result, the diffusion decoder and the deterministic repair decoder operate on different encoder layers. The deterministic-diffusion shared and deterministic-specific embeddings, along with skip connections, are processed separately by each decoder to obtain the final score and deterministic features:
$\mathbf{est_{deter}}&=\text{Dec}_{\text{deter}}(\mathbf{e^{\prime}_{dd}},t)\\ \mathbf{est_{score}}&=\text{Dec}_{\text{score}}(\mathbf{e^{\prime}_{s}},\mathbf{e^{\prime}_{dd}},\mathbf{e^{\prime}_{ds}},t)\end{aligned}\right.$ (23)
where $\mathbf{est_{score}}$ and $\mathbf{est_{deter}}$ represent the score for diffusion and the repaired deterministic feature, respectively. The training loss of the network is the sum of the score error and the complex spectrogram-based MSE:
@@
To further enhance RDM-SE, we propose utilizing both deterministic-only and deterministic-noisy features as conditions. Fig. 2 shows the structure of the dual-streaming encoding RDM-SE (DERDM-SE). Different from RDM-SE, both deterministic-noisy and deterministic-only features are used as conditions.
@@
where $\text{Enc}^{a}_{\text{DD}}$ and $\text{Enc}^{b}_{\text{DD}}$ represent the deterministic-diffusion shared layers with deterministic-only and deterministic-noisy conditions. Another T-RF bin Processing Block is used to fuse $\mathbf{e^{\prime}_{da}}$ and $\mathbf{e^{\prime}_{db}}$ as $\mathbf{e^{\prime}_{dd}}$. The diffusion-specific layers, bottleneck processing module, and deterministic information-repaired decoder follow the same processing as described in Equation (21), Equation (22), and Equation (26). The score estimation decoder is processed as follow:
@@
IV Experiments
IV-A Evaluation Datasets
IV-B Settings for the Baseline Models
  1. -
  2. -
  3. -
    TF-GridNet: The embedding dimension was 128, and the number of hidden units in the LSTM was 200. The self-attention mechanism had 4 heads. The approximate dimension of the frame-level key and value tensors was 512. The kernel size for unfolding and DeConv1D was 3, with the hop size of 1. The model included 4 GridNetBlocks. The STFT parameters included the window length of 512, the hop size of 128, and the FFT length of 512, using the Hann window function.
All baseline diffusion models used the same Noise Conditional Score Network (NCSN++) architecture (the number of parameters was 65.6 million (M)) as SGMSE+ [17]. The real and imaginary parts of the complex spectrograms were used as input feature. Only the deterministic enhanced features were used as the condition for diffusion, referred to as Deter-Diffusion. We compared StoRM (the 65.6M NCSN++ architecture was also used). Different from the original paper [25], the deterministic models for StoRM in this study used different structures instead of NCSN++. For GP-Unified [19], all encoder and decoder settings were the same as SGMSE+, resulting in 106M training parameters and 65.6M decoding parameters. All diffusion models were trained for 150 epochs.
TABLE I: Performance of Deter-Diffusion, StoRM, GP-Unified, and the proposed RDM-SE with DCCRN, DPRNN, and TF-GridNet enhanced conditions on the evaluation sets of the CHiME4 dataset. $\text{D}_{c}$ Decoder represents the use of the deterministic repair decoder.
Systems #T.P. #D.P. Simulated Real
Csig Cbak Covl SDR W-P STOI dMOS Csig Cbak Covl dMOS uMOS
Mixture - - 2.61 1.92 1.88 7.54 1.27 0.870 2.69 2.08 1.47 1.46 2.47 1.58
   + RDM-SE 25.2 25.2 3.58 2.75 2.88 11.65 2.22 0.950 3.61 3.43 4.00 3.12 3.67 3.46
     + Dual-Stream 26.1 24.0 3.48 2.74 2.81 11.50 2.17 0.949 3.66 3.40 4.08 3.13 3.73 3.49
   + RDM-SE 24.2 24.2 3.64 2.82 2.96 11.97 2.32 0.948 3.64 3.43 3.92 3.09 3.62 3.43
     + Dual-Stream 25.1 23.0 3.56 2.80 2.90 11.82 2.28 0.948 3.66 3.40 4.03 3.12 3.69 3.53
   + RDM-SE 30.9 30.9 3.60 2.82 2.93 12.30 2.28 0.951 3.63 3.19 4.01 2.90 3.47 2.95
     + Dual-Stream 31.8 29.7 3.50 2.82 2.84 12.40 2.21 0.952 3.70 3.21 4.09 2.95 3.58 3.09
IV-C Settings for the Proposed Methods
For the proposed deterministic method, the STFT parameters included the window length of 512, the hop size of 128, and the FFT length of 512, with the Hann window function. The settings for the encoder and decoder layers were the same as those in DCCRN. The embedding dimension within the T-E, T-RF, and T-F processing blocks was 48. The input and output dimensions of the intra- and inter-LSTM within the T-E and T-RF processing blocks were 144 and 96, respectively. The input and output dimensions of the intra- and inter-LSTM within the T-E processing blocks were 768 and 48, respectively. Additional specific settings are provided in the experimental section. For the proposed RDM-SE model, 3 deterministic-diffusion shared encoder layers with 1 T-RF processing block were used; 5 diffusion-specific encoder layers were used; 3 deterministic decoder layers and 8 diffusion decoder layers were used. For the proposed DERDM-SE model, we reduced the number of channels in the deterministic-diffusion shared encoder layers by half to ensure that the performance improvement was not due to an increase in the number of parameters. We have released the source code for the proposed model at: https://github.com/hshi-speech/Repair-Diffusion.
IV-D Additional Settings for Evaluation
In the diffusion process, we used PC sampling with 30 steps. In previous work [20], we found that within a 30-step reverse diffusion process, the sampling diversity at step 20 (with $t$ progressing from $T$ to 00) can be maintained as if starting from the initial state. Thus, in this paper, we start the diffusion process at $t=20$ instead of $t=30$. In previous work [20], we found that when the ensemble number was greater than 8, the performance difference became negligible. Therefore, this paper used the ensemble with a maximum of 8 samples.
IV-E Evaluation Metrics
Multiple evaluation metrics were used to analyze the improvements in human listening experience. For the simulated data, signal distortion (Csig) [37], background intrusiveness (Cbak) [37], overall quality (Covl) [37], the wideband perceptual evaluation of speech quality (WB-PESQ) [38], the short-time objective intelligibility (STOI) [39], and signal to distortion ratio (SDR) were used as evaluation metrics. For the real data, the subjective Mean Opinion Score (dMOS) was used, which was evaluated with the open-source toolkit DNSMOS [40, 41]. UTMOS (uMOS) [42] was also used.
IV-F Evaluation of Different Diffusion Models
Table I shows the performance of different deterministic models in combination with various diffusion models. Compared to the noisy mixture signal, SGMSE+ effectively enhances noisy signals. Deter-Diffusion, which used only deterministic enhanced features as conditions, performed better than SGMSE+. When using DCCRN-enhanced features as conditions, the performance of objective evaluation metrics (Csig, Cbak, Covl, SDR, WB-PESQ, STOI) on the simulated dataset was comparable to that of SGMSE+. This may be due to the signal distortions introduced by DCCRN. The DCCRN-Deter-Diffusion performed better at noise reduction, achieving the higher Cbak value, while the noisy SGMSE+ method excelled at speech signal recovery, yielding the higher Csig value. Although DCCRN-Deter-Diffusion did not outperform SGMSE+ in simulated objective evaluation metrics, it showed certain advantages in MOS: Deter-Diffusion achieved higher scores on both dMOS and uMOS on the real evaluation sets. The quality of the deterministic enhanced features affected the final performance of Deter-Diffusion. The performance ranking of the deterministic models on the simulated evaluation sets: TF-GridNet $>$ DPRNN $>$ DCCRN. However, TF-GridNet-Deter-Diffusion performed the worst on the simulated data. While the performance on the real evaluation sets, was inconsistent with that on the simulated evaluation sets: DPRNN $>$ TF-GridNet $>$ DCCRN. TF-GridNet outperformed DPRNN in dMOS but underperformed in uMOS. The TF-GridNet-Deter-Diffusion and DPRNN-Deter-Diffusion performed worse in dMOS on the real evaluation sets than the DCCRN-Deter-Diffusion.
In Table I, we also compared two models that incorporate the original noisy features as conditions: StoRM and GP-Unified. The models with deterministic-noisy condition demonstrated greater sensitivity to the deterministic enhanced features, regardless of whether deterministic information was incorporated into the diffusion model implicitly or explicitly during training.
GP-Unified explicitly incorporates deterministic information repair by using a deterministic decoder during training. Compared to StoRM, GP-Unified had comparable performance on the simulated evaluation sets, while had better performance on the real evaluation sets. For all dMOS and uMOS values on the real evaluation sets, GP-Unified outperformed StoRM. However, compared to Deter-Diffusion, GP-Unified still showed a decline in some metrics. Only with TF-GridNet-enhanced conditions, GP-Unified achieved better performance on the simulated evaluation sets (WB-PESQ, STOI). For the real evaluation sets, GP-Unified still had some disadvantages in dMOS values, which were similar to those of StoRM: with DCCRN or DPRNN conditions, both StoRM and GP-Unified had a decline in dMOS values. By comparing Deter-Diffusion, StoRM, and GP-Unified across various metrics on the real evaluation sets, we found that significant declines in noise reduction (Cbak) for StoRM and GP-Unified. This may explain the decrease in dMOS values: [43] indicates that the uMOS is insensitive to speech with noise compared to dMOS. For uMOS values, with DCCRN or DPRNN conditions, GP-Unified outperformed Deter-Diffusion. However, with TF-GridNet conditions, although it showed some improvement over StoRM, it still had significant performance degradation compared to Deter-Diffusion. Based on the performance of TF-GridNet as the condition on simulated and real evaluation sets, we argue that the T-F bin-level fine-grained processing, while achieving best performance on simulated data, may introduce distortions more easily to real data due to distribution differences between real and simulated data. However, the advantages brought by fine-level processing cannot be ignored for deterministic models. Besides, although DCCRN has lower dMOS and uMOS values than TF-GridNet in real evaluations sets, its performance was acceptable and demonstrates more stable diffusion performance, particularly in models that combine deterministic-noisy features.
Table I shows our proposed RDM-SE. We also used DCCRN, DPRNN, and TF-GridNet as deterministic models for diffusion. Compared to StoRM and GP-Unified, the proposed RDM-SE achieved better and more stable performance. On the real evaluation sets, RDM-SE outperformed StoRM and GP-Unified in almost all deterministic models. Compared to StoRM and GP-Unified, RDM-SE achieved better noise suppression, particularly on the real evaluation sets, with higher Cbak values. RDM-SE showed a greater advantage in uMOS compared to StoRM and GP-Unified: using all deterministic models, RDM-SE achieved consistent improvements compared to other deterministic-noisy diffusion models. For dMOS, different deterministic models had varying impacts on RDM-SE: with DCCRN and DPRNN, RDM-SE outperformed StoRM and GP-Unified; while with TF-GridNet, RDM-SE had comparable performance to StoRM and GP-Unified. Compared to Deter-Diffusion, RDM-SE showed significant improvements when using DCCRN, comparable performance when using DPRNN, and a slight decline when using TF-GridNet. Surprisingly, RDM-SE with DCCRN outperformed those using DPRNN and TF-GridNet, achieving the best performance. This did not occur in StoRM or GP-Unified. As previously analyzed, fine-grained processing may result in more severe distortion issues when handling real data. This suggests that while the proposed RDM-SE can mitigate distortion issues to some extent, its performance is still closely tied to the input deterministic feature. Explicitly incorporating deterministic repair ($+\text{D}_{c}$ Decoder) further improved the model’s performance, especially when TF-GridNet was used as the condition. In addition, given the performance differences among different deterministic methods under deterministic-only conditions (Deter-Diffusion) and deterministic-noisy conditions (StoRM and GP-Unified), we proposed to use dual-stream encoding (+ Dual-Stream). Although the dual-stream encoding model some performance degradation on simulated metrics (Csig, Cbak, Covl, SDR, WB-PESQ, STOI) compared to RDM-SE ($+\text{D}_{c}$ Decoder), it achieved improvements in dMOS and uMOS, leading to a better human hearing experience.
@@ IV-F Evaluation of Different Diffusion Models
TABLE II: Performance of the proposed COFFEE in the Evaluation Sets of CHiME4 dataset.
ID. T-E T-RF T-F #Params (M) Macs (G/s) Csig $\uparrow$ Cbak $\uparrow$ Covl $\uparrow$ SDR (dB) $\uparrow$ WB- PESQ $\uparrow$ STOI $\uparrow$
@@
TABLE III: Performance of the proposed deterministic COFFEE for different diffusion models on the evaluation sets of the CHiME4 dataset. $\text{D}_{c}$ Decoder represents the use of the deterministic repair decoder.
Systems #T.P. #D.P. Simulated Real
Csig Cbak Covl SDR W-P STOI dMOS Csig Cbak Covl dMOS uMOS

### v2

@@ III-A Deterministic Method: COFFEE
where $\mathbf{Y}_{or}$ and $\mathbf{Y}_{oi}$ are the real and imaginary parts of $\mathbf{Y}_{o}$, respectively. The mean $\mathbb{E}[\cdot]$ is computed over the batch dimension. The matrix $V^{-1/2}\in\mathbb{R}^{2\times 2}$ denotes the inverse square root of the $2\times 2$ covariance matrix constructed from the real and imaginary components. The learnable affine transformation is represented by $W^{a}$, and the bias vector by $\mathbf{b}_{r}$ and $\mathbf{b}_{i}$. Finally, $\operatorname{PReLU}(\cdot)$ denotes the element-wise parametric ReLU activation function applied separately to each component.
@@
where the $W^{b}$ is implemented with the T-E processing block. $B$ is the embedding extracted by the encoder. Fig. 2 shows the structure of the T-E processing Block. It contains an Intra-LSTM, an Inter-LSTM, and an attention mechanism to enhance full-band and sub-band modeling capabilities. Two additional linear layers with complex convolution are included for separate modeling. The Intra-LSTM, Inter-LSTM, and Attention are similar to the process of processing features described in Section II-A. The T-E bin level feature is processed with a large dimension reduction compared with the original T-F spectrogram.
@@
Finally, after the feature is upsampled to the same resolution as the original noisy feature, a T-F processing Block is used to make the final refinement. The T-F processing block has the same structure as the T-RF processing block.
@@
$CRM=\frac{\mathbf{Y}_{r}\mathbf{X}_{0r}+\mathbf{Y}_{i}\mathbf{X}_{0i}}{\mathbf{Y}_{r}^{2}+\mathbf{Y}_{i}^{2}}+j\frac{\mathbf{Y}_{r}\mathbf{X}_{0i}-\mathbf{Y}_{i}\mathbf{X}_{0r}}{\mathbf{Y}_{r}^{2}+\mathbf{Y}_{i}^{2}}$ (19)
where $\mathbf{Y}_{r}$ and $\mathbf{Y}_{i}$ represent the real and imaginary parts of the noisy complex spectrogram $\mathbf{Y}$, respectively. The $\mathbf{X}_{0r}$ and $\mathbf{X}_{0i}$ represent the real and imaginary parts of the clean complex spectrogram $\mathbf{X}_{0}$, respectively. According to the polar coordinates of Cartesian coordinate representation $\tilde{M}=\tilde{M}_{r}+j\tilde{M}_{i}$:
$\left\{\begin{aligned} \tilde{M}_{\text{mag}}&=\sqrt{\tilde{M}_{r}^{2}+\tilde{M}_{i}^{2}},\\ \tilde{M}_{\text{phase}}&=\arctan 2(\tilde{M}_{i},\tilde{M}_{r}).\end{aligned}\right$. (20)
@@
$\tilde{\mathbf{X}}_{0}=\mathbf{Y}_{mag}\cdot\tilde{M}_{mag}\cdot e^{j*(\mathbf{Y}_{phase}+\tilde{M}_{phase})}$ (21)
where $\mathbf{Y}_{mag}$ and $\mathbf{Y}_{phase}$ represent the magnitude and phase information of the noisy speech. The training loss is based on the MSE between the clean and enhanced spectrogram, which is the same as in Eq. (5)
III-B Diffusion Model: D2U-SE
@@
$\mathbf{X^{\prime}}=\underset{\text{COFFEE}}{\text{Deterministic}}(\mathbf{Y})$ (22)
@@
Figure 3: Flowchart of the proposed D3U-SE. The skip connections between the encoder and decoder are omitted in the figure for clarity. Lowercase letters denote the time-domain representations of their corresponding uppercase symbols.
We propose the D2U-SE to better utilize deterministic information. It is based on a dual-decoder structure. The model contains an encoder, a bottleneck processing module, a diffusion decoder, and a deterministic decoder. Within the encoder, the feature is processed by several deterministic-diffusion shared layers and diffusion-specific layers. The conditions and diffused features are first processed by the encoder:
$\begin{cases}\mathbf{e^{\prime}_{dd}}=\text{Enc}_{\text{DD}}(\mathbf{X^{\prime}}_{t},\mathbf{X^{\prime}},\mathbf{Y},t)\\ \mathbf{e^{\prime}_{ds}}=\text{Enc}_{\text{DS}}(\mathbf{e^{\prime}_{dd}},t),\end{cases}$ (27)
where $\text{Enc}_{\text{DD}}$ and $\text{Enc}_{\text{DS}}$ represent the deterministic-diffusion shared layers and diffusion-specific layers, respectively. $\mathbf{e^{\prime}_{dd}}$ and $\mathbf{e^{\prime}_{ds}}$ denote the deterministic-diffusion shared embeddings and diffusion-specific embeddings, respectively. The deterministic-diffusion shared layer consists of several ResNet blocks and a T-RF bin processing block. It performs downsampling only along the frequency axis to preserve the deterministic information. The deterministic-specific layer consists of several ResNet blocks. The attention layer is also applied to some layers, but not all. It performs downsampling either along the time axis or the time-frequency axis.
@@
It should be noted that not all encoder layers were used for incorporating the deterministic information. As a result, the diffusion decoder and the deterministic decoder operate on different encoder layers. The deterministic-diffusion shared and deterministic-specific embeddings, along with skip connections, are processed separately by each decoder to obtain the final score and deterministic features:
$\left\{\begin{aligned} \mathbf{est_{deter}}&=\text{Dec}_{\text{deter}}(\mathbf{e^{\prime}_{dd}},t)\\ \mathbf{est_{score}}&=\text{Dec}_{\text{score}}(\mathbf{e^{\prime}_{s}},\mathbf{e^{\prime}_{dd}},\mathbf{e^{\prime}_{ds}},t)\end{aligned}\right$. (29)
where $\mathbf{est_{score}}$ and $\mathbf{est_{deter}}$ represent the score for diffusion and the estimated deterministic feature, respectively. The training loss of the network is the sum of the score error and the complex spectrogram-based MSE:
@@
To further enhance D2U-SE, we propose utilizing both deterministic-only and deterministic-noisy features as conditions. Fig. 3 shows the structure of the D3U-SE. Different from D2U-SE, both deterministic-noisy and deterministic-only features are used as conditions.
@@
where $\text{Enc}^{a}_{\text{DD}}$ and $\text{Enc}^{b}_{\text{DD}}$ represent the deterministic-diffusion shared layers with deterministic-only and deterministic-noisy conditions. Another T-RF bin Processing Block is used to fuse $\mathbf{e^{\prime}_{da}}$ and $\mathbf{e^{\prime}_{db}}$ as $\mathbf{e^{\prime}_{dd}}$. The diffusion-specific layers, bottleneck processing module, and deterministic decoder follow the same processing as described in Equation (27), Equation (28), and Equation (32). The score estimation decoder is processed as follows:
@@
TABLE I: Performance of the proposed COFFEE in the Evaluation Sets of the CHiME4 dataset.
ID. T-E T-RF T-F
#Params
Macs
(G/s)
Csig $\uparrow$ Cbak $\uparrow$ Covl $\uparrow$
SDR
(dB) $\uparrow$
W-P $\uparrow$ STOI $\uparrow$
@@
Mixture - - 2.61 1.92 1.88 7.54 1.27 0.870
DCCRN 3.7 5.6 3.58 2.66 2.82 11.12 2.10 0.939
DPRNN 2.7 85.5 3.73 2.76 2.98 11.30 2.27 0.942
TF-GridNet 9.4 153.1 3.97 2.95 3.19 12.57 2.41 0.957
IV Experiments
IV-A Evaluation Datasets
IV-B Settings for the Baseline Models
  1. -
  2. -
  3. -
    TF-GridNet: The embedding dimension was 128, and the number of hidden units in the LSTM was 200. The self-attention mechanism had 4 heads. The approximate dimensions of the frame-level key and value tensors were 512. The kernel size for unfolding and DeConv1D was 3, with the hop size of 1. The model included 4 GridNetBlocks. The STFT parameters included the window length of 512, the hop size of 128, and the FFT length of 512, using the Hann window function.
All baseline diffusion models used the same noise scenarioal Score Network (NCSN++) architecture (the number of parameters was 65.6 million (M)) as SGMSE+ [15]. The real and imaginary parts of the complex spectrograms were used as input features. Only the deterministic enhanced features were used as the condition for diffusion, referred to as Deter-Diffusion. We compared StoRM (the 65.6M NCSN++ architecture was also used). Different from the original paper [22], the deterministic models for StoRM in this study used different structures instead of NCSN++. For GP-Unified [17], all encoder and decoder settings were the same as SGMSE+, resulting in 106M training parameters and 65.6M decoding parameters. All diffusion models were trained for 150 epochs.
IV-C Settings for the Proposed Methods
For the proposed deterministic method, the STFT parameters included the window length of 512, the hop size of 128, and the FFT length of 512, with the Hann window function. The settings for the encoder and decoder layers were the same as those in DCCRN. The embedding dimension within the T-E, T-RF, and T-F processing blocks was 48. The input and output dimensions of the intra- and inter-LSTM within the T-E and T-RF processing blocks were 144 and 96, respectively. The input and output dimensions of the intra- and inter-LSTM within the T-E processing blocks were 768 and 48, respectively. Additional specific settings are provided in the experimental section. For the proposed D2U-SE model, 3 deterministic-diffusion shared encoder layers with 1 T-RF processing block were used; 5 diffusion-specific encoder layers were used; 3 deterministic decoder layers and 8 diffusion decoder layers were used. For the proposed D3U-SE model, we reduced the number of channels in the deterministic-diffusion shared encoder layers by half to ensure that the performance improvement was not due to an increase in the number of parameters. We have released the source code for the proposed model at: https://github.com/hshi-speech/Repair-Diffusion.
TABLE II: Performance of Deter-Diffusion, StoRM, GP-Unified, and the proposed RDM-SE with DCCRN, DPRNN, and TF-GridNet enhanced conditions on the evaluation sets of the CHiME4 dataset. $\text{D}_{c}$ Decoder represents the use of the deterministic decoder. #Parameters represents the number of parameters.
Systems #Parameters Simulated Real
Train Inference Csig Cbak Covl SDR W-P STOI dMOS Csig Cbak Covl dMOS uMOS
Mixture - - 2.61 1.92 1.88 7.54 1.27 0.870 2.69 2.08 1.47 1.46 2.47 1.58
   + D2U-SE 25.2 25.2 3.58 2.75 2.88 11.65 2.22 0.950 3.61 3.43 4.00 3.12 3.67 3.46
   + D3U-SE 26.1 24.0 3.48 2.74 2.81 11.50 2.17 0.949 3.66 3.40 4.08 3.13 3.73 3.49
   + D2U-SE 24.2 24.2 3.64 2.82 2.96 11.97 2.32 0.948 3.64 3.43 3.92 3.09 3.62 3.43
   + D3U-SE 25.1 23.0 3.56 2.80 2.90 11.82 2.28 0.948 3.66 3.40 4.03 3.12 3.69 3.53
   + D2U-SE 30.9 30.9 3.60 2.82 2.93 12.30 2.28 0.951 3.63 3.19 4.01 2.90 3.47 2.95
   + D3U-SE 31.8 29.7 3.50 2.82 2.84 12.40 2.21 0.952 3.70 3.21 4.09 2.95 3.58 3.09